OpenStudy (anonymous):

integral (1+x^2)dx=x(1+2y)dy

OpenStudy (klimenkov):

Is \(y\) the function of \(x\)?

OpenStudy (anonymous):

yes

OpenStudy (klimenkov):

Calculte 2 integrals: \(\int\frac{1+x^2}{x}dx\) \(\int(1+2y)dy\)

OpenStudy (anonymous):

help with that please

OpenStudy (aravindg):

did u calculate that @IcekiD ?

OpenStudy (anonymous):

can u pls solve it for me.

OpenStudy (klimenkov):

\(\ln|x|+\frac{x^2}{2}=y+y^2+C\)

OpenStudy (anonymous):

thanks a lot